Two particles each of mass 1 kg are placed at P and Q such that PO=OQ=L. The gravitational force experienced by another 1 kg mass placed at R where OR=L is ?

Grade:

2 Answers

Shubham Pathy
5 years ago
r= distance of 1kg mass from P =distance of mass 1kg from Q
 
r^2=l^2+l^2 
=>  r=l√2
Now total gravitational force =2F=2GMm/r^2
Here F is force due to one particle
 
 
=2×G×1×1/2l2
= G/ll2
 
Ankur Singh
4 years ago
F(1)= Gm^2/r^2
       =G/2l^2
 F(2)=Gm^2/r^2
        =G/2l^2
 F(net)=√F(1)^2+F(2)^2
            =√(G/2l^2)^2+(G/2l^2)^2
            =√(G^2/4l^4)+(G^2/4l^4)
            =G/l^2√2